Could you help me solve for the ordered pairs?
y=5^(3x+1)
y=25^(-x)

Answer by scott8148(6628)

y=25^(-x) __ y=(5^2)^(-x) __ y=5^(-2x)

substituting __ 5^(3x+1)=5^(-2x) __ 3x+1=-2x __ subtracting 3x __ 1=-5x __ dividing by -5 __ -1/5=x

substituting __ y=25^(1/5)